the clays saved $3000 for their vacation, this includes $900 for travel ad $80 per day for miscellaneous, write an inequality which represents the number of days X, the family can stay on vacation
Answer by Theo(13342)

budget is 3000.
cost is 900 plus 80 per day.

900 + 80x must be less than or equal to 3000.

alternatively, 3000 must be greater than or equal to 900 + 80x.

that leads to the equation 3000 >= 900 + 80x.

subtract 900 from both sides of this equation to get 3000 - 900 >= 80x

divide both sides of this equation by 80 to get (3000 - 900) / 80 >= x

solve for x to get x <= 26.25.

round down because you can't have a partial day, and your answer is that the number of days they can spend on vacation is less than or equal to 26.

if 26 days, the total cost is 900 + 80 * 26 = 2980.
if 27 days, the total cost is 2980 + 80 = 3060.

at 27 days they are over budget.
at 26 days they are under budget.

most number of days is 26.